OpenStudy (anonymous):

translate each problem into an equation. (1) Brenda drove three times as far as Jan. (2) Brenda drove 24 miles more than Jan. How far did Jan drive?

OpenStudy (anonymous):

Let b represent Brenda's total distance. J is Jan's total distance 3j=b j+24=b -72=-2b -72/-3 = 36. B=36, J=12 Jan drove 12 miles, Brenda drove 36 miles.

OpenStudy (anonymous):

1. 3j=b 2. j+24=b